# Asymptotic Velocity Domination

### v1

@@ 1.  Introduction
An important property of polarized Gowdy cosmologies is that they admit an infinite set of Dirac observables [15]. Dirac observables are thought to capture the intrinsic ‘gauge invariant’ content of a gravity theory, but are normally elusive. In Appendix B we construct a new one-parameter family of the form
${\cal O}(\theta)=\int_{0}^{2\pi}\!\frac{dx^{1}}{2\pi}\,q(x^{0},x^{1};\theta)={\cal O}_{0}(\theta)+O(\rho^{2}/\lambda_{{\textsc{n}}}^{4})\,,\quad{\rm Im}\theta\neq 0\,.$ (1.1)
@@
Here $\rho>0$ is a temporal function whose $\rho=0$ level set can be identified with the Big Bang. Further, $(x^{0},x^{1})$ are the non-Killing coordinates of the Gowdy spacetime and $\lambda_{{\textsc{n}}}>0$ is the dimensionless reduced Newton constant. These quantities strongly Poisson commute with the constraints without using equations of motion. The leading term ${\cal O}_{0}(\theta)$ is in one-to-one correspondence to a family of Dirac observables in the VD system, while the subleading terms are organized alternatively according to powers of $\rho^{2}/\lambda_{{\textsc{n}}}^{2}$ or by the number of dynamical spatial derivatives. The integrand $q$ is linear in the Gowdy scalar $\phi$ (defining $M={\rm diag}(e^{\phi},e^{-\phi})$ in (2.1) below), and in a gauge fixed reduced phase space formulation only $\phi$ is canonically quantized, while the field $\tilde{\sigma}$ in (2.1) is treated as a renormalized composite operator in $\phi$. The gauge specializations single out coordinates $(\tau,\zeta)\in\mathbb{R}^{2}$, such that the Big Bang is located at $\tau\rightarrow$-$\infty$. In this framework ${\cal O}(\theta)$ becomes an on-shell conserved charge, $\partial_{\tau}{\cal O}^{\rm on}(\theta)=0$, with $q(\tau,\zeta+2\pi;\theta)=q(\tau,\zeta;\theta)$ spatially periodic and finite as $\tau\rightarrow$-$\infty$. The two-point function of the on-shell $q$ can be taken as the central object in the quantum theory, $\langle 0_{T}|q(\tau,\zeta;\theta)q(\tau^{\prime},\zeta^{\prime};\theta)|0_{T}\rangle$, where $|0_{T}\rangle$ is some (non-unique) Fock vacuum defining a ‘state’. Only its symmetric part is state dependent and on account of the linearity in $\phi$ its properties are coded by those of $\langle 0_{T}|\phi(\tau,\zeta)\phi(\tau^{\prime},\zeta^{\prime})+\phi(\tau^{\prime},\zeta^{\prime})\phi(\tau,\zeta)|0_{T}\rangle$ and its $\partial_{\tau},\partial_{\tau^{\prime}},\partial_{\tau}\partial_{\tau^{\prime}}$ derivatives. These are conveniently combined into a $2\times 2$ matrix two-point function ${\cal W}^{s}$. As described in Appendix A, one can view the velocity dominated Gowdy system as a Carroll-type gravity theory in its own right and follow analogous steps through VD Dirac observables and their integrand’s two-point function to find its properties coded by a matrix two-point function $\mathfrak{W}^{s}$ in the VD system analogous to ${\cal W}^{s}$. The question of quantum AVD can then be posed as follows: does ${\cal W}^{s}$ have a $\tau,\tau^{\prime}\rightarrow$-$\infty$ asymptotics that relates to its VD counterpart $\mathfrak{W}^{s}$? If so, can ${\cal W}^{s}$ be reconstructed from $\mathfrak{W}^{s}$? We find that the answers are affirmative for a large class of states we dub “time consistent”. Specifically, we obtain the
@@
$\quad+\sum_{l\geq 1}\sum_{k=0}^{l}e^{2k\tau}e^{2(l-k)\tau^{\prime}}\int\!\!\frac{d\zeta}{2\pi}\frac{d\zeta^{\prime}}{2\pi}f(\zeta)g(\zeta^{\prime})(\partial_{\zeta}\partial_{\zeta^{\prime}})^{l}I_{k}\,\mathfrak{W}^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})\,I_{l-k}^{T}\,.$ (1.2)
@@ 2.  Polarized Gowdy cosmologies and AVD
Gowdy cosmologies are spacetimes with two commuting spacelike Killing vectors $K_{1},K_{2}$. In adapted coordinates $X=(x^{0},x^{1},y^{1},y^{2})$ the line element can be brought into the form
$g^{\rm 2K}_{IJ}(X)dX^{I}dX^{J}=\gamma_{\mu\nu}(x)dx^{\mu}dx^{\nu}+\rho(x)M_{ab}(x)dy^{a}dy^{b}\,$, (2.1)
@@
i.e. $\gamma_{00}=e^{\tilde{\sigma}}(-n^{2}+s^{2})$, $\gamma_{01}=e^{\tilde{\sigma}}s=\gamma_{10}$, $\gamma_{11}=e^{\tilde{\sigma}}$. By a slight abuse of terminology we shall refer to $n,s$ as lapse, shift, respectively.111 For a two-dimensional Lorentzian metric the lapse $N$, shift $N^{1}$, spatial metric $\gamma_{11}$ proper would be $N=e^{\tilde{\sigma}/2}n$, $N^{1}=s$, $\gamma_{11}=e^{\tilde{\sigma}}$. Note that $e^{\tilde{\sigma}},n$ are spatial densities of weight 2,12,-1, respectively, and that $s$, being one-dimensional spatial vector, can also be viewed as a spatial density of weight 1-1. Cosmological spacetimes arise in this setting if $\rho$ is assumed to have everywhere timelike gradient, $\gamma^{\mu\nu}(x)\partial_{\mu}\rho\partial_{\nu}\rho<0$. This evaluates to $e_{0}(\rho)^{2}>n^{2}(\partial_{1}\rho)^{2}$, where $e_{0}(\rho)=(\partial_{0}-s\partial_{1})\rho$. The solutions generically have a Big Bang singularity, which in the coordinates singled out by $s=0,n=1,\rho=t$ occurs at the $t\rightarrow 0^{+}$ hypersurface. The evolution equations can be coded by a Lax pair and some of the integrable systems methodology can be adapted to this situation. In particular, there is a rich set of gravitational soliton solutions [7].
@@ 2.1 Action, proper time gauge, and reduced phase space
Here, $\lambda_{{\textsc{n}}}>0$ is the dimensionless reduced Newton constant, $e_{0}=\partial_{0}-{\cal L}_{s}$ (with ${\cal L}_{s}$ the one-dimensional Lie-derivative acting on spatial tensor densities) is the derivative transversal to the leaves of the $x^{0}={\rm const}.$ foliation, and $\sigma=\tilde{\sigma}+\frac{1}{2}\ln\rho$. We use $x^{1}\in\mathbb{R}$ but all fields in (2.3) are assumed to be spatially periodic with period $2\pi$ (and thus having a well-defined projection onto $S^{1}=\mathbb{R}/2\pi\mathbb{Z}$.) This form of the action is tailored towards holding the fields $\rho,\sigma,\phi$ fixed at the boundaries $x^{0}=t_{i},t_{f}$. Its Hamiltonian version (A.1) allows one to identify the canonically induced gauge symmetries (A.4). The subset (A.4,a,b) with the velocity-momentum relations inserted on the right hand side of (A.4b) yields the gauge variations $\delta_{\epsilon}^{\textsc{l}}$ of (2.3). Variation of $S^{\textsc{l}}$ with respect to $n,s$ gives the Lagrangian constraints
@@
$F_{0}(\rho,\phi)$ $\!\!=\!\!$ $\frac{e_{0}(\rho)}{e_{0}(\rho)^{2}\!-\!(n\partial_{1}\rho)^{2}}\Big{(}2n^{2}\partial_{1}^{2}\rho+\frac{\rho}{2}[(\partial_{0}\phi)^{2}+(n\partial_{1}\phi)^{2}]\Big{)}$
@@
- $\frac{n\partial_{1}\rho}{e_{0}(\rho)^{2}\!-\!(n\partial_{1}\rho)^{2}}\Big{(}2n^{2}\partial_{1}^{2}\rho+\frac{\rho}{2}[(\partial_{0}\phi)^{2}+(n\partial_{1}\phi)^{2}]\Big{)}\,.$
By assumption $e_{0}(\rho)^{2}\!-\!(n\partial_{1}\rho)^{2}>0$ for Gowdy cosmologies. Since $\partial_{1}e_{0}(\sigma+2\ln n)=e_{0}\partial_{1}(\sigma+2\ln n)$ the integrability condition $\partial_{1}[F_{0}+2e_{0}(\ln n)]=e_{0}[F_{1}/n+2\partial_{1}\ln n]$ must be satisfied. This is indeed the case subject to the evolution equations for $\rho,\phi$, and re-expresses the preservation (A.6) of the constraints under time evolution. The two first order relations can therefore locally be integrated to yield $\sigma$ as a functional of the on-shell $\rho$ and $\phi$, i.e. $\sigma=\sigma[\rho^{\rm on},\phi^{\rm on}]$. Further, $e_{0}[e_{0}(\sigma)-F_{0}]-\partial_{1}[\partial_{1}\sigma-F_{1}]=0$ reproduces the equation of motion for $\sigma$, i.e. $\delta S^{\textsc{l}}/\delta\rho=0$. That is, the functional $\sigma=\sigma[\rho^{\rm on},\phi^{\rm on}]$ is automatically on-shell as well. Moreover, if $\rho^{\rm on},\phi^{\rm on}$ are parameterized by their boundary values at $x^{0}=t_{i},t_{f}$, so will be $\sigma[\rho^{\rm on},\phi^{\rm on}]$, rendering $\sigma|_{t_{i}},\sigma|_{t_{f}}$ a functional of $\rho|_{t_{i}},\rho|_{t_{f}},\phi|_{t_{i}},\phi|_{t_{f}}$. Together, this allows one to take the fields $\rho,\phi$ as basic and to treat $\sigma$ as a composite field in them.
@@
where $F_{0}^{\rm pt}$, $F_{1}^{\rm pt}$ are obtained from $F_{0},F_{1}$ in (2.1) by the specialization to (2.6), (2.7) and the replacements $2n^{2}\partial_{1}^{2}\rho\mapsto 2\partial_{\zeta}^{2}\rho$, $2n^{2}\partial_{1}(e_{0}(\rho)/n)\mapsto 2\partial_{\zeta}\partial_{0}\rho$. The only vestige of the gravitational origin of the system then is that in order for $\sigma$ to be spatially periodic, one must have
@@
Here the counterpart of (2.9) needs to be imposed in order ensure that $\varsigma$ is spatially periodic.
@@ 2.2 The AVD property
$\varphi(t,\zeta)=\omega(\zeta)\ln(t/t_{0})+\varphi_{0}(\zeta)\,,\quad\varsigma(t,\zeta)=\ln(t/t_{0})\frac{1}{2}\omega(\zeta)^{2}+\varsigma_{0}(\zeta)\,$,
$\varsigma_{0}(\zeta)=-2\ln(\frac{n(\zeta)}{n(\zeta_{0})})+\int_{\zeta_{0}}^{\zeta}\!d\zeta^{\prime}\omega(\zeta^{\prime})\partial_{\zeta^{\prime}}\varphi_{0}(\zeta^{\prime})\,.$ (2.14)
@@
The origin of the exceptional points can be understood from the blow-up behavior of the curvature scalars. Considering ${\cal R}_{2}:=R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$, the result of [3], Thm. IV.1, is that the rate of blow up for $t\rightarrow 0^{+}$ is determined solely by the coefficient $\omega(\zeta)$ in the velocity dominated solution (LABEL:VDsol). Specifically,
${\cal R}_{2}(t,\zeta_{0})=\left\{\begin{array}[]{cc}O(t^{-3-\omega(\zeta_{0})^{2}})&\mbox{if}\;\;\omega(\zeta_{0})^{2}\neq 1\,,\\[5.69054pt] O(t^{-2})&\mbox{if}\;\;\omega(\zeta_{0})^{2}=1\,,\;(\partial_{\zeta}\omega)(\zeta_{0})\neq 0\,,\\[5.69054pt] O(\ln^{2}t/t_{0})&\mbox{if}\;\;\omega(\zeta_{0})^{2}=1\,,\;(\partial_{\zeta}\omega)(\zeta_{0})=0\,,\;\;(\partial_{\zeta}^{2}\omega)(\zeta_{0})\neq 0\,.\end{array}\right$. (2.19)

### v2

@@ 1.  Introduction
An important property of polarized Gowdy cosmologies is that they admit an infinite set of Dirac observables [15]. Dirac observables are thought to capture the intrinsic ‘gauge invariant’ content of a gravity theory, but are normally elusive, see e.g. [16]. In Appendix B we construct a new one-parameter family of the form
${\cal O}(\theta)=\int_{0}^{2\pi}\!\frac{dx^{1}}{2\pi}\,\jmath_{0}^{{\textsc{p}}}(x^{0},x^{1};\theta)={\cal O}_{0}(\theta)+O(\rho^{2}/\lambda_{{\textsc{n}}}^{2})\,,\quad{\rm Im}\theta\neq 0\,.$ (1.1)
@@
Here $\rho>0$ is a temporal function whose $\rho=0$ level set can be identified with the Big Bang. Further, $(x^{0},x^{1})$ are the non-Killing coordinates of the Gowdy spacetime and $\lambda_{{\textsc{n}}}>0$ is the dimensionless reduced Newton constant. These quantities strongly Poisson commute with the constraints without using equations of motion. The leading term ${\cal O}_{0}(\theta)$ is in one-to-one correspondence to a family of Dirac observables in the VD system, while the subleading terms are organized alternatively according to powers of $\rho^{2}/\lambda_{{\textsc{n}}}^{2}$ or by the number of dynamical spatial derivatives. The integrand $\jmath_{0}^{{\textsc{p}}}$ is linear in the Gowdy scalar $\phi$ (defining $M={\rm diag}(e^{\phi},e^{-\phi})$ in (2.1) below) and its momentum $\pi^{\phi}$, and in a gauge fixed reduced phase space formulation only $(\phi,\pi^{\phi})$ is canonically quantized, while the field $\tilde{\sigma}$ in (2.1) is treated as a renormalized composite operator. Suitable gauge specializations single out coordinates $(\tau,\zeta)\in\mathbb{R}^{2}$, such that the Big Bang is located at $\tau\rightarrow$-$\infty$. Then ${\cal O}(\theta)$ becomes an on-shell conserved charge, $\partial_{\tau}{\cal O}^{\rm on}(\theta)=0$, whose integrand $q(\tau,\zeta;\theta)$ differs from $\jmath_{0}^{{\textsc{p}}}(e^{\tau},\zeta;\theta)$ by a total $\partial_{\zeta}$ derivative, and stays finite as $\tau\rightarrow$-$\infty$. The two-point function of the on-shell $q$ can be taken as the central object in the quantum theory, $\langle 0_{T}|q(\tau,\zeta;\theta)q(\tau^{\prime},\zeta^{\prime};\theta)|0_{T}\rangle$, where $|0_{T}\rangle$ is some (non-unique) Fock vacuum defining a ‘state’. Only its symmetric part is state dependent and on account of the linearity in $\phi$ its properties are coded by those of $\langle 0_{T}|\phi(\tau,\zeta)\phi(\tau^{\prime},\zeta^{\prime})+\phi(\tau^{\prime},\zeta^{\prime})\phi(\tau,\zeta)|0_{T}\rangle$ and its $\partial_{\tau},\partial_{\tau^{\prime}},\partial_{\tau}\partial_{\tau^{\prime}}$ derivatives. These are conveniently combined into a $2\times 2$ matrix two-point function ${\cal W}^{s}$. As described in Appendix A, one can view the velocity dominated Gowdy system as a Carroll-type gravity theory in its own right and follow analogous steps through VD Dirac observables and their integrand’s two-point function to find its properties coded by a matrix two-point function $\mathfrak{W}^{s}$ in the VD system analogous to ${\cal W}^{s}$. The question of quantum AVD can then be posed as follows: does ${\cal W}^{s}$ have a $\tau,\tau^{\prime}\rightarrow$-$\infty$ asymptotics that relates to its VD counterpart $\mathfrak{W}^{s}$? If so, can ${\cal W}^{s}$ be reconstructed from $\mathfrak{W}^{s}$? We find that the answers are affirmative for a large class of states we dub “time consistent”. Specifically, we obtain the
@@
$\qquad+\sum_{l\geq 1}\sum_{k=0}^{l}e^{2k\tau}e^{2(l-k)\tau^{\prime}}\int\!\!\frac{d\zeta}{2\pi}\frac{d\zeta^{\prime}}{2\pi}f(\zeta)g(\zeta^{\prime})(\partial_{\zeta}\partial_{\zeta^{\prime}})^{l}I_{k}\,\mathfrak{W}^{s}(\tau,\tau^{\prime},\zeta\!-\!\zeta^{\prime})\,I_{l-k}^{T}\,.$ (1.2)
@@ 2.  Polarized Gowdy cosmologies and AVD
Gowdy cosmologies are exact solutions of the 4-dimensional Einstein equations with two commuting spacelike Killing vectors $K_{1},K_{2}$. In adapted coordinates $X=(x^{0},x^{1},y^{1},y^{2})$ the line element can be brought into the form
$g^{\rm 2K}_{IJ}(X)dX^{I}dX^{J}=\gamma_{\mu\nu}(x)dx^{\mu}dx^{\nu}+\rho(x)M_{ab}(x)dy^{a}dy^{b}\,.$ (2.1)
@@
i.e. $\gamma_{00}=e^{\tilde{\sigma}}(-n^{2}+s^{2})$, $\gamma_{01}=e^{\tilde{\sigma}}s=\gamma_{10}$, $\gamma_{11}=e^{\tilde{\sigma}}$. By a slight abuse of terminology we shall refer to $n,s$ as lapse, shift, respectively. [^1] For a two-dimensional Lorentzian metric the lapse $N$, shift $N^{1}$, spatial metric $\gamma_{11}$ proper would be $N=e^{\tilde{\sigma}/2}n$, $N^{1}=s$, $\gamma_{11}=e^{\tilde{\sigma}}$. Note that $e^{\tilde{\sigma}},n$ are spatial densities of weight 2,12,-1, respectively, and that $s$, being one-dimensional spatial vector, can also be viewed as a spatial density of weight 1-1. Cosmological spacetimes arise in this setting if $\rho$ is assumed to have everywhere timelike gradient, $\gamma^{\mu\nu}(x)\partial_{\mu}\rho\partial_{\nu}\rho<0$. This evaluates to $e_{0}(\rho)^{2}>n^{2}(\partial_{1}\rho)^{2}$, where $e_{0}(\rho)=(\partial_{0}-s\partial_{1})\rho$. The field equations are presented in (LABEL:evolpt) below. The solutions generically have a Big Bang singularity, which in the coordinates singled out by $s=0,n=1,\rho=t$ occurs at the $t\rightarrow 0^{+}$ hypersurface. The evolution equations can be coded by a Lax pair and some of the integrable systems methodology can be adapted to this situation. In particular, there is a rich set of gravitational soliton solutions [7].
@@ 2.1 Action, proper time gauge, and reduced phase space
This action can be inferred from [19, 13, 15] or directly verified using computer algebra. Here, $\lambda_{{\textsc{n}}}>0$ is the dimensionless reduced Newton constant, $e_{0}=\partial_{0}-{\cal L}_{s}$ (with ${\cal L}_{s}$ the one-dimensional Lie-derivative acting on spatial tensor densities) is the derivative transversal to the leaves of the $x^{0}={\rm const}.$ foliation, and $\sigma=\tilde{\sigma}+\frac{1}{2}\ln\rho$. We use $x^{1}\in\mathbb{R}$ but all fields in (2.3) are assumed to be spatially periodic with period $2\pi$ (and thus having a well-defined projection onto $S^{1}=\mathbb{R}/2\pi\mathbb{Z}$.) This form of the action is tailored towards holding the fields $\rho,\sigma,\phi$ fixed at the boundaries $x^{0}=t_{i},t_{f}$. Its Hamiltonian version (A.1) allows one to identify the canonically induced gauge symmetries (A.4). The subset (A.4,a,b) with the velocity-momentum relations inserted on the right hand side of (A.4b) yields the gauge variations $\delta_{\epsilon}^{\textsc{l}}$ of (2.3). Variation of $S^{\textsc{l}}$ with respect to $n,s$ gives the Lagrangian constraints
@@
$F_{0}(\rho,\phi)$ $\!\!=\!\!$ $\frac{e_{0}(\rho)}{e_{0}(\rho)^{2}\!-\!(n\partial_{1}\rho)^{2}}\Big(2n^{2}\partial_{1}^{2}\rho+\frac{\rho}{2}[e_{0}(\phi)^{2}+(n\partial_{1}\phi)^{2}]\Big)$
@@
- $\frac{n\partial_{1}\rho}{e_{0}(\rho)^{2}\!-\!(n\partial_{1}\rho)^{2}}\Big(2n^{2}\partial_{1}^{2}\rho+\frac{\rho}{2}[e_{0}(\phi)^{2}+(n\partial_{1}\phi)^{2}]\Big)\,.$
The fact that the constraints in 2-Killing vector reductions can be solved for one of the fields is well-known [7, 17]; we present (2.1) mostly to highlight that this can be done without gauge fixing or coordinate choices. By assumption, $e_{0}(\rho)^{2}\!-\!(n\partial_{1}\rho)^{2}>0$ for Gowdy cosmologies. Since $\partial_{1}e_{0}(\sigma+2\ln n)=e_{0}\partial_{1}(\sigma+2\ln n)$ the integrability condition $\partial_{1}[F_{0}+2e_{0}(\ln n)]=e_{0}[F_{1}/n+2\partial_{1}\ln n]$ must be satisfied. This is indeed the case subject to the evolution equations for $\rho,\phi$, and re-expresses the preservation (A.6) of the constraints under time evolution. The two first order relations can therefore locally be integrated to yield $\sigma$ as a functional of the on-shell $\rho$ and $\phi$, i.e. $\sigma=\sigma[\rho^{\rm on},\phi^{\rm on}]$. Further, $e_{0}(n^{-1}e_{0}(\sigma))-\partial_{1}(n\partial_{1}\sigma)=e_{0}(n^{-1}F_{0})-\partial_{1}F_{1}=2\partial_{1}^{2}n-e_{0}(\phi)^{2}/(2n)+n(\partial_{1}\phi)^{2}/2$ reproduces the equation of motion for $\sigma$, i.e. $\delta S^{\textsc{l}}/\delta\rho=0$. That is, the functional $\sigma=\sigma[\rho^{\rm on},\phi^{\rm on}]$ is automatically on-shell as well. Moreover, if $\rho^{\rm on},\phi^{\rm on}$ are parameterized by their boundary values at $x^{0}=t_{i},t_{f}$, so will be $\sigma[\rho^{\rm on},\phi^{\rm on}]$, rendering $\sigma|_{t_{i}},\sigma|_{t_{f}}$ a functional of $\rho|_{t_{i}},\rho|_{t_{f}},\phi|_{t_{i}},\phi|_{t_{f}}$. Together, this allows one to take the fields $\rho,\phi$ as basic and to treat $\sigma$ as a composite field in them.
@@
where $F_{0}^{\rm pt}$, $F_{1}^{\rm pt}$ are obtained from $F_{0},F_{1}$ in (2.1) by the specialization to (2.6), (2.7). Explicitly,
$F_{0}^{\rm pt}(\rho,\sigma)$ $\!\!=\!\!$ $\frac{\partial_{0}\rho}{(\partial_{0}\rho)^{2}\!-\!(\partial_{\zeta}\rho)^{2}}\Big(2\partial_{\zeta}^{2}\rho+\frac{\rho}{2}[(\partial_{0}\phi)^{2}+(\partial_{\zeta}\phi)^{2}]\Big)$
- $\frac{\partial_{\zeta}\rho}{(\partial_{0}\rho)^{2}\!-\!(\partial_{\zeta}\rho)^{2}}\big(2\partial_{\zeta}\partial_{0}\rho+\rho\partial_{0}\phi\partial_{\zeta}\phi\big)$
$F_{1}^{\rm pt}(\rho,\sigma)$ $\!\!=\!\!$ $\frac{\partial_{0}\rho}{(\partial_{0}\rho)^{2}\!-\!(\partial_{\zeta}\rho)^{2}}\big(2\partial_{\zeta}\partial_{0}\rho+\rho\partial_{0}\phi\partial_{\zeta}\phi\big)$ (2.9)
- $\frac{\partial_{\zeta}\rho}{(\partial_{0}\rho)^{2}\!-\!(\partial_{\zeta}\rho)^{2}}\Big(2\partial_{\zeta}^{2}\rho+\frac{\rho}{2}[(\partial_{0}\phi)^{2}+(\partial_{\zeta}\phi)^{2}]\Big)\,.$
The integrability condition mentioned after Eq.(2.5) specializes to $\partial_{\zeta}F_{0}^{\rm pt}=\partial_{0}F_{1}^{\rm pt}$, and can be checked to hold on account of the first two equations in (2.8). Note that the $\phi$ field equation is through $\rho$ explicitly $(x_{0},\zeta)$-dependent; in particular, it is not conformally invariant. The other vestige of the gravitational origin of the system then is that in order for $\sigma$ to be spatially periodic, one must have
@@
Here the counterpart of (2.10) needs to be imposed in order to ensure that $\varsigma$ is spatially periodic.
@@ 2.2 The AVD property
$\varphi(t,\zeta)=v(\zeta)\ln(t/t_{0})+\varphi_{0}(\zeta)\,,\quad\varsigma(t,\zeta)=\ln(t/t_{0})\frac{1}{2}v(\zeta)^{2}+\varsigma_{0}(\zeta)\,$,
$\varsigma_{0}(\zeta)=-2\ln(\frac{n(\zeta)}{n(\zeta_{0})})+\int_{\zeta_{0}}^{\zeta}\!d\zeta^{\prime}v(\zeta^{\prime})\partial_{\zeta^{\prime}}\varphi_{0}(\zeta^{\prime})\,.$ (2.15)
@@
The origin of the exceptional points can be understood from the blow-up behavior of the curvature scalars. Considering ${\cal R}_{2}:=R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$, the result of [3], Thm. IV.1, is that the rate of blow up for $t\rightarrow 0^{+}$ is determined solely by the coefficient $v(\zeta)$ in the velocity dominated solution (LABEL:VDsol). Specifically,
${\cal R}_{2}(t,\zeta_{0})=\left\{\begin{array}[]{cc}O(t^{-3-v(\zeta_{0})^{2}})&\mbox{if}\;\;v(\zeta_{0})^{2}\neq 1\,,\\[5.69054pt] O(t^{-2})&\mbox{if}\;\;v(\zeta_{0})^{2}=1\,,\;(\partial_{\zeta}v)(\zeta_{0})\neq 0\,,\\[5.69054pt] O(\ln^{2}t/t_{0})&\mbox{if}\;\;v(\zeta_{0})^{2}=1\,,\;(\partial_{\zeta}v)(\zeta_{0})=0\,,\;\;(\partial_{\zeta}^{2}v)(\zeta_{0})\neq 0\,.\end{array}\right$. (2.20)